# Finding needles in a haystack: A Black-Box Approach to Invisible Watermark Detection

### v1

@@ 1 $\mathcal{D}_{dt}\leftarrow\mathcal{D}_{d}$ for $e=1$ to $E$ do
             $p_{c},p_{dt}\leftarrow f(\mathcal{B}_{c};\theta),f(\mathcal{B}_{dt};\theta)$   // Get model outputs
             $\mathcal{L}_{total}\leftarrow\mathcal{L}_{sm}(p_{c})+\mathcal{L}_{lin}(p_{dt})$   // Calculate the losses
             $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{total}$   // Update the model with gradients
3
@@
             $(p_{d})\text{Ranked}\leftarrow rank(f(\mathcal{D}_{d};\theta))$   // Get model outputs and rank
@@
             $\theta\leftarrow init(\theta)$   // Reinitialize model to avoid overfitting
6
7
@@ 4.2 Watermark Detection
Table 2: Watermark detection performance of WMD across different datasets and watermark methods. Methods marked with $*{}^{*}$ represent generative watermarks, which are directly embedded during the image generation process. Higher AUC and TPR @ 0.1 FPR indicate better performance, while lower FPR @ 0.9 TPR is desirable.
@@ 4.3 Ablation Study
Minimize
@@ Appendix 0.A Detailed Experiments Setting & Hyperparameters
Watermark Embedded Bits Parameters Values
@@ 0.B.1 Watermark removal attacks
HiDDeN SS SSig TR
@@ 0.B.2 Filtering invisible anomalies
Data Poisoning Dataset Tracing Unlernable Example
@@ 0.B.3 Synthesis clean data
0% 20% 40% 60% 80%
@@ Appendix 0.C Further Ablation Study
Clean

### v3

@@ 3.4 Overall Method
             $p_{c},p_{dt}\leftarrow f(\mathcal{B}_{c};\theta),f(\mathcal{B}_{dt};\theta)$
              // Get model outputs
             $\mathcal{L}_{total}\leftarrow\mathcal{L}_{sm}(p_{c})+\mathcal{L}_{lin}(p_{dt})$
              // Calculate the losses
             $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{total}$
              // Update the model with gradients
6
@@
8
            $(p_{d})\text{Ranked}\leftarrow rank(f(\mathcal{D}_{d};\theta))$
              // Get model outputs and rank
@@
             $\theta\leftarrow init(\theta)$
              // Reinitialize model to avoid overfitting
10
11
@@ 4.2 Watermark Detection
Table 2: Watermark detection performance of WMD across different datasets and watermark methods. Methods marked with represent generative watermarks, which are directly embedded during the image generation process. Higher AUC and TPR @ 0.1 FPR indicate better performance, while lower FPR @ 0.9 TPR is desirable.
@@ 4.3 Ablation Study
{tblr}
cells = c, cell13 = c=4, cell31 = r=4, vline3 = 1, vline3-6 = 2,4-6, vline2-6 = 3, vline2-6 = 4, vline2-6 = 5, vline2-6 = 6, hline1,3,7 = -, hline2 = 3-6, hline4-6 = 2-6, & Minimize
@@ Appendix 0.A Detailed Experiments Setting & Hyperparameters
{tblr}
cell31 = r=2, cell32 = r=2, cell51 = r=4, cell52 = r=4, cell91 = r=6, cell92 = r=6, cell161 = r=2, cell162 = r=2, hline2-3,5,9,15-16, 18, hline2 = 1.3pt, Watermark & Embedded Bits Parameters Values
@@ 0.B.1 Watermark removal attacks
{tblr}
cells = c, vline2 = -, hline2 = -, & HiDDeN SS SSig TR
@@ 0.B.2 Filtering invisible anomalies
{tblr}
colspec = X[c]X[c]X[c]X[c]X[c], cells = c, cell12 = c=2, cell15 = c=2, vline2-5 = 1, vline2-6 = 2-5, hline1,3-6 = -, hline2 = 2-6, & Data Poisoning Dataset Tracing Unlernable Example
@@ 0.B.3 Synthesis clean data
{tblr}
cells = c, cell22 = r=4, vline2-6 = 1-2, vline2,3-6 = 3-5, hline1-2,6 = -, hline3-5 = 1,3-6, & 0% 20% 40% 60% 80%
@@ Appendix 0.C Further Ablation Study
{tblr}
roweven = c, row1 = c, row5 = c, cell13 = c=6, cell31 = r=3, cell32 = c, cell33 = c, cell34 = c, cell35 = c, cell36 = c, cell37 = c, cell38 = c, vline3 = 1, vline3-8 = 2,4-5, vline2-8 = 3, vline2-8 = 4, vline2-8 = 5, hline1-3,6 = -, hline4-5 = 2-8, & Clean